You would like to slow down your aging process by
Answer #1

given

T = 26.2 years = 8.262 X 108 sec

t = 18.9 years = 5.960 X 108 sec

and we have C = 3 X 108 m/s

using the formula

T = t / ( \sqrt{}(1 - v2 / c2 ) )

( 8.262 X 108 )2 = ( 5.960 X 108 ) 2 / ( (1 - v2 / (3 X 108)2 ) )

(1 - v2 / (3 X 108)2 ) = 0.520

v2 / (3 X 108)2 = 0.48

v2 = 4.32 X 1016

v = 2.078 X 108 m/sec

the required speed is v = 2.078 X 108 m/sec